OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

its first one.....

OpenStudy (anonymous):

the equation of the line in slope-intercept form is y=mx+c -3=(-6)(-2)+c -3=12+c -15=c c=-15 so the equation is y=-6x-15 m is the slope